OpenStudy (anonymous):

Find the Taylor series for f(x) centered at the given value of a. f(x)=ln(x) a=8. Find the associated radius of convergence R.

OpenStudy (anonymous):

\[\ln (8)- \sum_{n=1}^{\infty}\frac{ (\frac{ -1 }{ 8 })^2(x-8)^n }{ n }\]

OpenStudy (anonymous):

THat's the correct Taylor series series but how do I find the radius of convergence?

OpenStudy (anonymous):

Never mind. I got it.